# Veri-Car: Towards Open-world Vehicle Information Retrieval

### v3

@@ 5.5 Model Configurations
In order to properly asses our model and choice of loss function we benchmark our hierarchical MS loss against different versions of the loss and its non-hierarchical baseline. We tested the following configurations:
  • Cross-Entropy: We replace the linear layers with a linear classifier for both make-type-mode-year and color heads.
  • MS: Original MS loss (Equation 1).
@@
  • HiSupCon: Original version of the loss as proposed by Zhang et al. [45], as seen in Equation 2. Original SupCon [46] is used to train the color head.
@@
  • HiMS-Max: Uses the top-down approach, as HiSupCon, but with MS loss as the base objective.
@@
  • HiMS-Min (Ours): Employs the bottom-up approach to the label hierarchy, using MS loss as the base objective, as described in Equation 3.
All configurations that use a variation of MS as the loss for the make-type-mode-year head, use the original MS formulation to train the color head.
We will also benchmark our proposed Combined model against the Standalone Make, Type, Model, Year and Color models found in the Appendix.
5.6 Training details
@@
The Stanford Cars dataset includes ground truth bounding box locations of the cars in the images, so we crop the car with a 5% margin to ensure we do not leave important features out. After cropping, we resize the image to $256\times 256$. Then, to increase the variety of samples, and prevent overfitting, we randomly crop an area of the image of at least 50% and resize it to $224\times 224$. Finally, we randomly flip the image horizontally. All configurations were trained with an Adam optimizer with a cyclic learning rate scheduling. We have selected to use the Adam optimizer as it is still the standard algorithm for training deep learning models at this scale, due to its fast convergence speed. The scheduler follows a triangular linear cycling schedule due to the fact that models tend to need some help escaping local minimums. Pushing the learning rate up tends to help to not only converge faster, but reach better local minimums on the loss surface [47]. Our scheduler cycles through a minimum learning rate of $2\times 10^{-6}$ and maximum of $2\times 10^{-4}$ for 200 epochs. We set the batch size to 600 for Cross-Entropy, MS, HiMS-Min and HiMS-Max and 350 for HiSupCon, due to memory restrictions. All experiments were run on 4 A10 GPUs.
@@ 5.7 Evaluation Metrics
During inference, the Make, Type, Model, Year and Color Retrieval model uses $K$-NN to assign a label to a sample embedding as illustrated in Figure 4. Embeddings for Make, Type, Model, Year Retrieval and Color Retrieval are extracted independently by their corresponding embedding head. Distances are calculated using euclidean distance.
@@ 5.8 Results
The upcoming section lays-out the training details, presents and analyzes the results of the experiments on Make, Type, Model, Year retrieval and color retrieval.
5.8.1 Make, Type, Model, Year Retrieval
Table 2(a) shows the retrieval performance for Make, Type, Model, Year labels when only seen classes exist in the database. Results show that in this case HiMS-Min outperforms baselines on both Precision@1 and mAP@R. Table 2(b) shows similar results on a combined dataset that takes into account both seen and unseen classes. Results in Table 2(b) show that HiMS-Min produces better separation of seen and unseen classes than other methods. Tables 2(c) and 2(d) show the breakdown of the combined performance found in Table 2(b). Here we can see that HiMS-Min beats most methods in both metrics in the unseen set, except HiMS-Max which beats it in Precision@1, while beating baselines on Precision@1 in the seen set. This points towards HiMS-Min being able to find general features that distinguish between car models even if they were not seen during training. We believe these effects are caused by creating relationships at all levels of the hierarchy, but not inflating the importance at any one level. General results in all tables in Table 2 show that training using HiMS-Min produces higher quality embeddings with more predictive capabilities. And this can be observed not only over baselines not using hierarchies at all, but over the ones that use the ”top-down” approach shown in Equation 2. Furthermore, Figure 5(a) shows the progressive fall in the performance of the model as we add unseen classes to the database and test on them. The performance trend of HiMS-Min and HiMS-Max point not only towards HiMS-Min being more robust to classes being added into the database, but models with hierarchical supervision being able to accommodate more classes into the same embedding space.
@@ 5.8.1 Make, Type, Model, Year Retrieval
Method mAP@R Precision@1 HiSupCon 52.05 90.86 Cross-Entropy 76.17 93.95 MS 94.07 95.71 HiMS-Max 93.03 95.44 HiMS-Min 94.13 95.92
@@
Method mAP@R Precision@1 HiSupCon 48.09 89.36 Cross-Entropy 62.18 91.41 MS 83.90 93.56 HiMS-Max 83.51 93.65 HiMS-Min 84.37 94.11
@@
Method mAP@R Precision@1 HiSupCon 48.78 89.36 Cross-Entropy 69.88 92.61 MS 90.96 94.44 HiMS-Max 90.18 94.23 HiMS-Min 90.94 94.82
@@
Method mAP@R Precision@1 HiSupCon 44.91 88.06 Cross-Entropy 27.15 85.94 MS 51.17 89.56 HiMS-Max 53.17 90.99 HiMS-Min 54.42 90.86
@@
To properly asses the quality of the model, we study the minimum amount of samples per-class needed in the database to correctly classify an image from an unseen class. Therefore, we subsample our unseen classes, add them into the seen database, and then test retrieval of our unseen query set. Figure 5(b) shows the average result over 100 runs. We can observe that using hierarchical supervision helps MS trained models to be robust to having less samples in the database, with HiMS-Min scoring $81.6\%$ and HiMS-Max scoring $82.88\%$ with 12 samples in the database. We believe this is due to the hard positive and negative mining employed by MS loss. Learning only through these samples at all levels reduces the bias introduced by easy, or redundant samples, helping the model focus on the distinctive features at all levels of hierarchy.
As explained in Section 5.4, the model should be able to correctly assign labels at higher levels of the hierarchy to the samples it incorrectly predicted at the last level of the hierarchy. Table 3 shows that as we go up in the hierarchy, models get better at classifying samples that were incorrectly classified at the Year level (lowest level) of the hierarchy. Additionally, we can see that HiMS-Max and HiSupCon outperform HiMS-Min. This was expected due to the lesser importance placed on creating relationships at higher hierarchical levels. Nonetheless, HiMS-Min still outperforms MS at the Make level of the hierarchy. Along with that, all models performed poorly at the Make-Type-Model level due to the dataset not having a diversity of years associated to each model.
Table 4 shows the retrieval comparison between our combined model and a standalone Make, Type, Model and Year model. In general, we can see that our model performs comparatively to the benchmark across most settings in both embedding quality and classification ability except on the unseen split. Table 4(d) shows that we lose both embedding quality and classification ability on the unseen set with a loss of 7.88% and 2.22% respectively. More details on the standalone Make, Type, Model and Year model can be found on the Appendix C.1.
At last, due to nature of the application, a robustness to weather conditions assessment is necessary. We created a subset of the Stanford dataset, which we dubbed ”Bad Weather”, which contains 61 images where it is raining, snowing, cloudy and/or the camera lens was stained. Images were collected from our test set, samples can be found in Section B. Results found in Table 5 show that our proposed method (HiMS-Min) is quite robust to this changes in the environment, only losing 1.58% and 4.64% in Precision@1 and mAP@R respectively.
@@
Method Precision@1 Make-Type-Model Make-Type Make HiSupCon 1.80 46.72 81.31 Cross-Entropy 1.24 41.09 69.06 MS 1.05 48.25 75.87 HiMS-Max 1.32 53.29 81.91 HiMS-Min 2.20 46.15 76.56
@@
Table 4: Retrieval performance comparison between our Make, Type, Model, Year and Color Combined model and Make, Type, Model and Year Standalone Model.
Method mAP@R Precision@1 Combined Model 94.16 95.92 Standalone Model 94.11 96.18
(a) Seen
Method mAP@R Precision@1 Combined Model 84.37 94.11 Standalone Model 86.02 94.78
(b) Combined
Method mAP@R Precision@1 Combined Model 90.94 94.82 Standalone Model 91.16 95.15
(c) Seen with combined database
Method mAP@R Precision@1 Combined Model 54.42 90.86 Standalone Model 62.30 93.08
(d) Unseen with combined database
Table 5: Bad weather subset make, type, model and year retrieval results.
Test Set mAP@R Precision@1 Full Test set 94.13 95.92 Bad Weather 89.70 94.34
5.8.2 Color Retrieval
Table 6(a) reports the Color Retrieval performance on our automatically generated labels for the Stanford Cars dataset. Although we do not beat MS and HiMS-Max in color classification, we perform well and are consistent across the tested performance metrics with Precision1 reaching 88.78% and 87.65% on the seen and unseen set respectively, having better classification power than Cross-Entropy and HiSupCon baselines. Furthermore, results show that using the hierarchy to train the make-type-model-year head affects the capacity of the model to learn embeddings for color. We believe this is due to the extra supervision placed on the rest of the characteristics over color. Nevertheless, it appears to disproportionately affect HiMS-Min over HiMS-Max. We believe this due to the fact that a model that was pre-trained on such a large amount of data using contrastive learning already has a hierarchical representation close to the one we produce using HiMS-Min. Therefore, the supervision provided by such a loss function only reinforces the same structure while HiMS-Max gives it the opportunity to reconfigure in a way that allows for color embeddings to be more properly optimized. More evidence is the superior accuracy of HiMS-Max at higher hierarchical levels seen in Table 3.
Figure 6 shows the confusion matrix for the Predicted versus True color labels. The errors of the model are reasonable since most of the mistakes happen between colors that are similar or that under specific lighting conditions they might look similar or the same. Specifically, note the combinations of pink and red, gold and yellow, as well as purple and brown. At the same time, in each of those combinations there is one color that belongs to a class which is severely underrepresented in our dataset (Table 1).
Additionally, we want to study how much prediction performance is lost due to the auto-generated labels and using a combined model. In this experiment we trained our combined model as we stated in Section 5.6 and trained two standalone models, one on Stanford Cars and the other on VCoR’s training set. More details on the latter two models can be found in the Appendix C.2. All models are tested on VCoR’s test set since it is the only manually labeled dataset we have at hand. Table 7 shows that our combined model trained on our auto-generated labels keeps 89.92% of the classification ability of a standalone color model. Additionally, our combined model only loses 1.68% against the standalone version trained on Stanford Cars. This points towards the label distribution and induced noise due to the labeling process to be the culprit of the performance drop.
Finally, as in the previous section, we have to test for weather robustness. Results in Table 8 show that our model learned color features that are sensitive to changes in weather, experiencing a 9.54% and 13% loss in Precision@1 and mAP@R respectively. This is to be expected since features like snow affect how colors are reflected. As it was shown by Goodfellow et al. [49], even changes that we deem small can constitute great shifts in what a neural network perceives, thus rendering them ineffective.
Table 6: (a) Seen set retrieval-based color classification results. (b) Unseen set retrieval-based color classification results. Both are trained and evaluated using the Stanford Cars dataset.
Method Color mAP@R Precision@1 Top-1 Acc. HiSupCon 6.83 58.67 Cross-Entropy 35.90 86.22 88.39 MS 88.13 88.79 HiMS-Max 87.78 88.69 HiMS-Min 87.01 88.78
(a) Seen
Method Color mAP@R Precision@1 Top-1 Acc. HiSupCon 6.68 57.23 Cross-Entropy 34.53 84.31 87.99 MS 84.85 87.85 HiMS-Max 84.48 87.94 HiMS-Min 83.96 87.65
(b) Unseen
Table 7: VCoR color retrieval results.
Method Training Data Precision@1 Combined Stanford Cars 81.42 Standalone Stanford Cars 83.10 Standalone VCoR 90.55
Figure 6: HiMS-Min loss color confusion matrix. Each cell reports the percentage of observations falling in the combination of predicted vs true labels.
Table 8: Results of the bad weather subset color retrieval.
Test Set mAP@R Precision@1 Full Test set 87.01 88.78 Bad Weather 74.58 79.24
@@ 6 OOD Detection
OOD detection is a fundamental component of the system. We utilize KNN+ [42], an existing distance-based OOD detection method. KNN+ is a non-parametric approach that uses the distance of features in the feature space to their $k_{\text{th}}$ nearest neighbor to distinguish between ID and OOD samples. Level set estimation on this metric determines the threshold above which a new sample is considered OOD. The method can be fine-tuned by searching for the $k_{\text{th}}$ neighbor’s distance that achieves a 95% recall of ID samples while minimizing the false positive rate (FPR) on OOD samples. Figure 7 illustrates the inference flow of new samples within the OOD framework. This figure zooms in on the OOD detection module in Figure 1 in which car images are taken as input, and the output is whether the image is OOD or ID. Training set samples pass through the same pre-processing and Trained Encoder Network as for Make, Type, Model, Year and Color Retrieval, leveraging only the make-type-model-year head, to generate embedding vectors, thereby creating the ID manifold in the embedding space. Similarly, a new image/sample is processed through the same modules to generate its embedding vector, which is then mapped to the embedding space.
@@ 6.3 Results
Table 9 shows OOD results for models trained with hierarchical losses and MS loss. For KNN+ experiments, we ran an exhaustive search for the best $k$ and landed on $k=1$. Sun et al. [42] showed that datasets with a small number of samples per-class do not need to retrieve many to detect OOD samples. In our case, the Stanford Cars dataset averages 41 samples per class in the database, which is relatively small, and thus it makes sense that our chosen $k$ is lower than in other datasets. Results show that HiMS-Min outperforms all other methods in both AUROC and FPR@95, using KNN+, by 1.51% on the latter. We believe this is tied to the capacity of HiMS-Min to better separate unseen classes from seen classes as seen by the higher mAP@R in Tables 2(d) and 2(b). Furthermore, all methods perform better when using KNN+, which could be due to the lack of a sufficient number samples in the ID set to properly create a Gaussian ellipsoid, as required by the Mahalanobis distance [50]. In addition, we test the performance of the OOD module while adding new classes into the database. Figure 8(a) shows the reduction in FPR@95 and AUROC performance, as we add new classes. In this experiment half of the classes were kept as unseen and the other half were added into the database one by one. We can observe that HiMS-Min is more robust to new classes being added than other methods. The results provided are averaged over 100 runs. The order in which the 36 available classes are added into the database is randomized for each of the runs.
Finally, Figure 8(b) shows the trade-off between new class recall and FPR@95 as we add samples from a new class into the database. Results shown are the average over 20 random runs for all 36 unseen classes. We can observe that all models tend to have a severe drop-off in recall the less samples of the new class are present in the database. As expected, it appears that the higher the FPR@95, the easier it is to recall samples from a new class. However, HiMS-Min appears to have a better trade-off than baselines since at all levels of N samples it is closer to the optimal trade-off symbolized by the star at the top-left corner. We believe this is tied to the AUROC score. We can observe on Table 9 that as AUROC increases, the better trade-off we have. Moreover, HiMS-Min only needs 12 samples in the database to recall 82.20% of the new class samples as ID, only a 13.48% loss w.r.t the 95% TPR we calibrated the OOD module for. At the same time, as we already observed in Figure 5(b), HiMS-Min scores 81.6% in Precision@1 a 10.19% loss w.r.t. the Precision@1 observed in Table 2(d). This results show that we only need to collect and label 12 samples per-new-class in order for HiMS-Min to have good predictive capabilities.
@@
Method KNN+ Mahalanobis FPR95 $\downarrow$ AUROC $\uparrow$ FPR95 $\downarrow$ AUROC $\uparrow$ HiSupCon 62.21 82.26 91.89 58.76 Cross-Entropy 46.34 89.69 75.47 77.49 MS 32.19 92.05 84.64 70.41 HiMS-Max 35.42 91.14 82.23 70.17 HiMS-Min 30.61 92.12 88.04 66.89
Figure 8: (a) Effect of adding samples of a new class in FPR@95 and AUROC (b) Recall vs FPR@95 as new class samples are added to the database.

### v4

@@ 4.3.3 Model Configurations
In order to properly asses our model and choice of loss function we benchmark our hierarchical multi-similarity loss against different versions of the loss and its non-hierarchical baseline. We tested the following configurations:
  • MS: Original Multi-similarity loss (Equation 1).
@@
  • HiSupCon: Original version of the loss as proposed by Zhang et al. [39], as seen in Equation 2.
@@
  • HiMS-Max: Uses the top-down approach, as HiSupCon, but with Multi-Similarity loss as the base objective.
@@
  • HiMS-Min (Ours): Employs the bottom-up approach to the label hierarchy, using Multi-Similarity loss as the base objective, as described in Equation 3.
4.3.4 Training details
@@
The Stanford Cars dataset includes ground truth bounding box locations of the cars in the images, so we crop the car with a 5% margin to ensure we do not leave important features out. After cropping, we resize the image to $256\times 256$. Then, to increase the variety of samples, and prevent overfitting, we randomly crop an area of the image of at least 50% and resize it to $224\times 224$. Finally, we randomly flip the image horizontally. All configurations are trained with an Adam optimizer with a cyclic learning rate scheduling. The scheduler has a minimum learning rate of $2\times 10^{-6}$ and maximum of $2\times 10^{-4}$, we cycle through learning rates as in  [43] for 200 epochs. We set the batch size to 600 for MS, HiMS-Min and HiMS-Max and 350 for HiSupCon, due to memory restrictions. All experiments were run on 4 A10 GPUs.
4.4 Color Retrieval
In the following subsection, we give the details on the training configuration, architecture and dataset used to train the Color Retrieval model.
4.4.1 Dataset
As mentioned in previous sections, the color retrieval task was trained independently of the rest of the system since we found no open source datasets that included all relevant car characteristics. For this reason, the color label cannot be included in the hierarchy, and the Color Retrieval model is thus trained on the Kaggle Vehicle Color Recognition (VCoR) [4] which includes 15 color classes and 10.5K images.
4.4.2 Architecture
Similar to the Make, Type, Model, Year retrieval model we leverage OpenCLIP’s ViT-B/16 [41] trained on LAION-2B [42] as the backbone to our architecture and add a 128 dimension linear head at the end to produce the embeddings.
4.4.3 Model Configurations
In order to properly asses our model and choice of loss function we benchmark Multi-Similarity loss against Cross Entropy and Rachmadi et al. [24] Two-Stream Cross Entropy approach:
  • MS: Original Multi-similarity loss (Equation 1) with a linear layer of size 128 placed on top of the backbone.
4.4.4 Training Details
As in the Make, Type, Model, Year Retrieval model, we resize the images to $256\times 256$, randomly crop an area of the image of at least 66%, and resize again it to $224\times 224$. Finally, we randomly flip the image horizontally with a probability of 50%. All models were trained with an Adam optimizer with a cyclic learning rate scheduling. The Multi-similarity model’s scheduler has minimum learning rate of $2\times 10^{-6}$ and maximum of $2\times 10^{-4}$ and cycles through learning rates as in [43] for 200 epochs. While the Cross Entropy model’s cycles through a minimum of $5\times 10^{-6}$ and maximum of $1\times 10^{-4}$ for 50 epochs.
@@ 4.5 Evaluation Metrics
During inference, the Make, Type, Model, Year and Color Retrieval models use $k$-NN to assign a label to a sample embedding as illustrated in Figure 3. Embeddings for Make, Type, Model, Year Retrieval and Color Retrieval are extracted independently by their corresponding model. Distances are calculated using euclidean distance.
@@ 4.6 Results
4.6.1 Make, Type, Model, Year Retrieval Model
Table 1(a) shows the retrieval performance for Make, Type, Model, Year Model when only seen classes exist in the database. Results show that in this case HiMS-Min outperforms baselines on Prec@1 while coming in second, behind Multi-similarity, on mAP@R. Table 1(b) shows similar results on a combined dataset that takes into account both seen and unseen classes. Results in Table 1(b) show that HiMS-Min and Multi-Similarity can produce better separation of seen and unseen classes than other methods. Tables 1(c) and 1(d) show the breakdown of the combined performance found in Table 1(b). Here we can see that HiMS-Min beats most methods in both metrics in the unseen set, except HiMS-Max which ties it in Prec@1, while beating baselines on Prec@1 in the seen set. This points towards HiMS-Min being able to find general features that distinguish between car models even if they were not seen during training. We believe these effects are caused by creating relationships at all levels of the hierarchy, but not inflating the importance at any one level. Furthermore, Figure 4(a) shows the progressive fall in the performance of the model as we add unseen classes to the database and test on them. The performance trend of HiMS-Min and Multi-similarity point towards HiMS-Min being more robust to classes being added into the database.
@@ 4.6.1 Make, Type, Model, Year Retrieval Model
Method mAP@R Prec@1 HiSupCon 68.24 93.20 Multi-Similarity 94.78 96.13 HiMS-Max 91.65 96.03 HiMS-Min 94.11 96.18
@@
Method mAP@R Prec@1 HiSupCon 64.30 91.83 Multi-Similarity 86.38 94.46 HiMS-Max 83.60 94.52 HiMS-Min 86.02 94.78
@@
Method mAP@R Prec@1 HiSupCon 65.18 91.77 Multi-Similarity 92.03 95.11 HiMS-Max 88.73 94.83 HiMS-Min 91.16 95.15
@@
Method mAP@R Prec@1 HiSupCon 60.23 92.11 Multi-Similarity 60.23 91.41 HiMS-Max 59.92 93.08 HiMS-Min 62.30 93.08
@@
To properly asses the quality of the model, we study the minimum amount of samples per-class needed in the database to correctly classify an image from an unseen class. Therefore, we subsample our unseen classes, add them into the seen database, and then test retrieval of our unseen query set. Figure 4(b) shows the average result over 100 runs. We can observe that using hierarchical supervision helps Multi-similarity trained models to be less sensitive to having less samples in the database, with HiMS-Min scoring $82.99\%$ with 8 samples in the database. We believe this is due to the hard positive and negative mining employed by MS loss. Learning only through these samples at all levels reduces the bias introduced by easy, or redundant samples, helping the model focus on the distinctive features at all levels of hierarchy.
As explained in Section 4.2, the model should be able to correctly assign labels at higher levels of the hierarchy to the samples it incorrectly predicted at the last level of the hierarchy. Table 2 shows that as we go up in the hierarchy, models get better at classifying samples that were incorrectly classified at the Year level (lowest level) of the hierarchy. Additionally, we can see that HiMS-Max and HiSupCon outperform HiMS-Min. This was expected due to the lesser importance placed on creating relationships at higher hierarchical levels. Nonetheless, HiMS-Min still makes a small gain, at the Make level of the hierarchy, over Multi-Similarity loss. Along with that, all models performed poorly at the Model level due to the dataset not having a diversity of years associated to each model.
@@
Method Prec@1 Make-Type-Model Make-Type Make HiSupCon 1.10 48.02 78.63 Multi-Similarity 1.16 48.84 77.13 HiMS-Max 1.51 52.08 78.87 HiMS-Min 1.18 46.67 77.25
@@
4.6.2 Color Retrieval Model
Table 3: Color retrieval model results.
Method Color mAP@R Prec@1 Top-1 Acc. Multi-Similarity 89.68 90.55 Cross Entropy 85.31 88.09 90.29 Two-Stream 84.29 87.53 89.52
Figure 5: Multi-similarity loss color confusion matrix. Each cell reports the percentage of observations falling in the combination of predicted vs true labels.
@@ 5 OOD Detection
Out-of-Distribution (OOD) detection is a fundamental component of the system. We utilize KNN+ [45], an existing distance-based OOD detection method. KNN+ is a non-parametric approach that uses the distance of features in the feature space to their $k_{\text{th}}$ nearest neighbor to distinguish between in-distribution (ID) and OOD samples. Level set estimation on this metric determines the threshold above which a new sample is considered OOD. The method can be fine-tuned by searching for the $k_{\text{th}}$ neighbor’s distance that achieves a 95% recall of ID samples while minimizing the false positive rate (FPR) on OOD samples. Figure 6 illustrates the inference flow of new samples within the OOD framework. Training set samples pass through the same pre-processing as for Make, Type, Model, Year Retrieval and the Trained Encoder Network to generate embedding vectors, thereby creating the in-distribution manifold in the embedding space. Similarly, a new image/sample is processed through the same modules to generate its embedding vector, which is then mapped to the embedding space.
@@ 5.3 Results
Table 4 shows OOD results for models trained with hierarchical losses and Multi-similarity loss. For KNN+ experiments, we ran an exhaustive search for the best $k$ and landed on $k=1$. We not only want to optimize the FPR, but we would also like to find the minimum amount of samples needed to be added to the ID database in order to recognize new class samples as ID. The search took into account both FPR@95 and the average recall when a single sample of a new class is added to the ID database. While $k=1$ met both objectives for all methods, $k=4$ achieves the lowest FPR@95 for HiMS-Min, and $k=1$ provided the best trade-off between average recall and FPR@95 (Figure 7). Sun et al. [45] showed that datasets with a small number of samples per-class do not need to retrieve many neighbours to detect OOD samples. In our case, the Stanford Cars dataset averages 41 samples per class in the database, which is relatively small, and thus it makes sense that our chosen $k$ is lower than in other datasets. All methods perform better when using KNN+, which could be due to the lack of a sufficient number samples in the ID set to properly create a Gaussian ellipsoid, as required by the Mahalanobis distance [46]. Moreover, Figure 7 shows that HiMS-Min outperforms all baselines in FPR@95 if we were to choose $k=4$. Experiments in Figure 8 not only shows a similar drop-off in performance when a single sample is added into the ID database, but a general decrease in performance as well for all methods. We believe this is due to, as the low mAP@R but high Prec@1 in Table 1(d) suggest, the quality of the clusters is not high leading to samples of unseen being more spread-out throughout the embedding space. Results shown are the average over 10 random runs for all 36 unseen classes. Furthermore, Figure 9(b) shows a simplified version of Figure 8. We can observe that HiSupCon only needs 4 samples to classify $>80\%$ of the samples of the new class as ID, while HiMS-Min and MS need 8 samples to achieve similar performance.
@@
Method KNN+ Mahalanobis FPR95 $\downarrow$ AUROC $\uparrow$ FPR95 $\downarrow$ AUROC $\uparrow$ HiSupCon 63.14 83.38 91.38 61.26 Multi-Similarity 28.44 92.98 83.89 72.32 HiMS-Max 35.18 91.23 81.79 73.97 HiMS-Min 28.72 93.10 85.47 70.80